[Hallc_running] [New Logentry] Hall Status and plan

ciprian at jlab.org ciprian at jlab.org
Wed Oct 13 08:40:03 EDT 2021

Logentry Text:
RC Meeting Information
# RC Daily Meetings are at 10 am, remote: https://bluejeans.com/861439920 (wiki page: https://hallcweb.jlab.org/wiki/index.php/RC_Meeting,_Wednesday,_October_13_,_2021)

Accelerator Status
# Beyond usual trips and machine hiccups we have received beam fairly smoothly, even with the hall A beamline being set up over night
# Setup for Hall A will continue on 1pass for at least 2 days
#* Their production data starts with 3pass so we will hear from them soon when they will want to do a pass change

Hall C Status
# The hall is hanging in there and experiment is in good shape
# So far we have collected approximately 760 pions in this configuration
# Good luck to Dave Gaskell in his double role as RC and spokesperson over the next week

Current Run Plan
# Continue data taking until early Nov

Current Issues
# The HB has seen some fluctuations in the LHe levels. At the next opportunistic access Heidi would like to investigate the cryo line
#* Heidi believes this is related to a vacuum issue and the access is for diagnostic purposes only. Can be escorted during regular office hours (do not call her in at midnight for this)
#* If indeed it proves to be the case the only remedy would be to warm up the magent and pump the region (which would be a multiple day operation)
#* So far the PID loops have done their job and Heidi (together with the shift crew) are keeping a close eye on the LHe levels


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3923396
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211013/6ce655d9/attachment.html>

More information about the Hallc_running mailing list